what is the measurement of height if the radius of the cylinder is 3cm and the volume is 121cm? Then graph this function.
This is what I got so far:
v=2pir^2 + 2pirh
121= 2pe(3)^2 + 2pi(3)h
121 = 18 + 6
h= 24
I don't know if this is right. Don't I need 2 points to plot anything on a graph? Could you please help me? Christine

Answer by fractalier(2101)

The volume of a right circular cylinder is found by
V = (pi)(r^2)h
If the radius is 3 and the volume is 121, we have
121 = (pi)(9)h and
h = 121 / 9(pi)